40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 20 16:14:17 2017, gautam, Update, Equipment loan, Impedance test kit borrowed from Downs 
    Reply  Thu Sep 21 01:32:00 2017, gautam, Update, ALS, AUX X Innolight AM measurement running Innolight_AM.pdfInnolight_AM.tar.gzIMG_7599.JPG20170921_203741_TFAG4395A_21-09-2017_115547_FourSquare.pdf
       Reply  Thu Sep 21 01:55:16 2017, rana, Update, ALS, X End table of Shame 20170921_003143.jpg20170921_002430.jpg20170921_002243.jpg20170921_001906.jpg
          Reply  Thu Sep 21 15:23:04 2017, gautam, Omnistructure, ALS, Long cable from LSC->IOO 
          Reply  Fri Sep 29 11:16:52 2017, Steve, Update, ALS, Y End table corrected ETMYf1.jpg
          Reply  Fri Oct 6 17:08:09 2017, Steve, Update, ALS, X End table beam traps corrected c1.jpgc2.jpgc3.jpgc4.jpg
       Reply  Sun Sep 24 20:47:15 2017, rana, Update, Computer Scripts / Programs, RF TF Uncertainties TFAG4395A_21-09-2017_115547_FourSquare.pdfTFAG4395A_21-09-2017_115547.tgz
          Reply  Wed Sep 27 00:20:19 2017, gautam, Update, ALS, More AM sweeps TFAG4395A_26-09-2017_202344_FourSquare.pdflowTemp.tgzTFAG4395A_26-09-2017_231630_FourSquare.pdfhighTemp.tgz
             Reply  Wed Sep 27 23:44:45 2017, gautam, Update, ALS, Proposed PM measurement setup IMG_7609.JPG
Message ID: 13337     Entry time: Wed Sep 27 23:44:45 2017     In reply to: 13335
Author: gautam 
Type: Update 
Category: ALS 
Subject: Proposed PM measurement setup 

Attachment #1 is a sketch of the proposed setup to measure the PM response of the EX NPRO. Previously, this measurement was done via PLL. In this approach, we will need to calibrate the DFD output into units of phase, in order to calibrate the transfer function measurement into rad/V. The idea is to repeat the same measurement technique used for the AM - take ~50 1 average measurements with the AG4395, and look at the statistics. 

Some more notes:

  • Delay line box is passive, just contains a length of cable.
  • IQ Demodulation is done using an aLIGO 1U chassis unit, with the actual demod board electronics being D0902745
  • The RF beatnote amplitude out of the IR beat PD is ~ -8dBm.
  • The ZHL-3A amplifiers have gain of 24dB, so the amplified beat should be ~16dBm
  • At the LSC rack, the amplified beat is split into two - one path goes to the LO input of D0902745 (so at most 13dBm), the other goes through the delay line.
  • On the demod board, the LO signal is amplified with a AP1053, rated at 10dB gain, max output of 26dBm, so the signal levels should be fine for us, even though the schematic says the nominal LO level is 10dBm - moreover, I've ignored cable losses, insertion losses etc so we should be well within spec.
  • The mixer is PE4140. The datasheet quotes LO levels of 17dBm for all the "nominal" tests, we should be within a couple of dBm of this number.
  • There is no maximum value specified for the RF input signal level to the mixer on the datasheet, but I expect it to be <10dBm.
  • We should park the beatnote around 30MHz as this should be well within the operational ranges for the various components in the signal chain.
Attachment 1: IMG_7609.JPG  1.284 MB  Uploaded Thu Sep 28 13:48:16 2017  | Hide | Hide all
IMG_7609.JPG
ELOG V3.1.3-